25.0 mL of 0.100 M HC2H3O2 is titrated by a 0.100 M NaOH. A) Write the net ionic equation. B) calculate the pH of the solution after adding 10 mL of NaOH


1
Expert's answer
2022-06-12T17:02:03-0400

Solution:

(A):

Molecular equation:

HC2H3O2(aq) + NaOH(aq) → NaC2H3O2(aq) + H2O(l)

Total ionic equation:

H+(aq) + C2H3O2(aq) + Na+(aq) + OH(aq) → Na+(aq) + C2H3O2(aq) + H2O(l)

Net ionic equation:

H+(aq) + OH(aq) → H2O(l)


(B):

Moles of HC2H3O2 = Molarity of HC2H3O2 × Volume of HC2H3O2

Moles of HC2H3O2 = (0.100 M) × (25.0 mL) × (1 L / 1000 mL) = 0.0025 mol

Moles of HC2H3O2 = 0.0025 mol

Moles of NaOH = Molarity of NaOH × Volume of NaOH

Moles of NaOH = (0.100 M) × (10 mL) × (1 L / 1000 mL) = 0.001 mol

Moles of NaOH = 0.001 mol


Balanced chemical equation:

HC2H3O2(aq) + NaOH(aq) → NaC2H3O2(aq) + H2O(l)


ICE table for the reaction:



According to the ICE table:

n(HC2H3O2) = 0.0015 mol

n(NaC2H3O2) = 0.0010 mol


Total volume of solution = Volume of HC2H3O2 + Volume of NaOH

Total volume of solution = (25.0 mL + 10.0 mL) × (1 L / 1000 mL) = 0.035 L

Molarity of solute = Moles of solute / Liters of solution

Therefore,

[HC2H3O2] = (0.0015 mol) / (0.035 L) = 0.04286 M

[NaC2H3O2] = (0.0010 mol) / (0.035 L) = 0.02857 M


The Henderson-Hasselbalch equation is commonly used to calculate the pH of a buffer solution from the concentration of the buffer components:



where pKa is the acid dissociation constant and [base] and [acid] are the base and acid concentrations, respectively, in the chemical equation.

In this specific reaction, the base is NaC2H3O2 and the acid is HC2H3O2

Ka for acetic acid (HC2H3O2) is 1.8×10−5

pKa = −log(Ka) = −log(1.8×10−5) = 4.745

Therefore,

pH = 4.745 + log(0.02857 M / 0.04286 M) = 4.569 = 4.57

pH = 4.57



Answer:

A) Net ionic equation: H+(aq) + OH(aq) → H2O(l)

B) pH = 4.57